# A Model Zoo on Phase Transitions in Neural Networks

### v1

@@ Abstract
Using the weights of trained Neural Network (NN) models as data modality has recently gained traction as a research field - dubbed Weight Space Learning (WSL). Multiple recent works propose WSL methods to analyze models, evaluate methods, or synthesize weights. Weight space learning methods require populations of trained models as datasets for development and evaluation. However, existing collections of models — called ‘model zoos’ — are unstructured or follow a rudimentary definition of diversity. In parallel, work rooted in statistical physics has identified phases and phase transitions in NN models. Models are homogeneous within the same phase but qualitatively differ from one phase to another. We combine the idea of ‘model zoos’ with phase information to create a controlled notion of diversity in populations. We introduce 12 large-scale zoos that systematically cover known phases and vary over model architecture, size, and datasets. These datasets cover different modalities, such as computer vision, natural language processing, and scientific ML. For every model, we compute loss landscape metrics and validate full coverage of the phases. With this dataset, we provide the community with a resource with a wide range of potential applications for WSL and beyond. Evidence suggests the loss landscape phase plays a role in applications such as model training, analysis, or sparsification. We demonstrate this in an exploratory study of the downstream methods like transfer learning or model weights averaging.
@@ 1 Introduction
There is growing research interest in addressing these gaps by analyzing collections of trained models. Using neural networks as data modality – dubbed weight space learning (WSL) has become its own research field [^1] ICLR 2025 hosts the first workshop on Neural Networks as Data Modality https://weight-space-learning.github.io/. Within the field of WSL, populations of models are used in a variety of domains: Applications include predicting how well models generalize (Jiang et al., 2020), or detecting poisoned models (IARPA, 2024). Other methods use WSL torwards Neural Architecture Search to identify the relation between weights and model performance (Martin et al., 2021; Zhou et al., 2024; Liu et al., 2024). Different work re-uses existing traiend weights to extend conventional transfer learning strategies (Yosinski et al., 2014; Mensink et al., 2021) by combining multiple models, via a combination of their outputs (Polikar, 2012; Ganaie et al., 2022; Mohammed and Kora, 2023), knowledge distillation (Liu et al., 2019; Shu et al., 2021), or by directly combining their weights (Wortsman et al., 2022a, b; Mitchell et al., 2022; Ainsworth et al., 2022; Ilharco et al., 2022; Rame et al., 2023; Ramé et al., 2024; Navon et al., 2024), to build combined models with improved in- and out-of-distribution performance.
@@
Another line of research investigates phase transitions in NNs and connects them to loss landscape metrics (Martin and Mahoney, 2019a; Yang et al., 2021). Motivated by statistical physics, they identify two main types of hyperparameters in NN training: the noisiness of the training process, dubbed temperature, and the amount of data relative to the size of the model, dubbed relative load. Using that notion, they identify distinct phases with qualitatively different model properties on the temperature-load landscape (Yang et al., 2021). Phase transitions have been studied extensively in the machine learning literature (Schwarze et al., 1992; Seung et al., 1992; Martin and Mahoney, 2019a, b). While phase transitions are typically studied theoretically under certain limits, e.g., an infinitely wide NN (Lewkowycz et al., 2020), we observe qualitatively similar behaviors for models with a practical size. Within phases, models are relatively homogeneous, with abrupt changes from one phase to the next. Along the load dimension, one example of a phase transition is double descent (Nakkiran et al., 2019). Neural scaling laws describe the load and temperature relation, e.g., as power laws to train models to specific phases (Hestness et al., 2017; Rosenfeld et al., 2020; Gordon et al., 2021; Hoffmann et al., 2022; Sorscher et al., 2022).
Figure 1: Overview of the approach. Left: Phases of Neural Networks with qualitatively different behavior have been identified (Martin and Mahoney, 2019a; Yang et al., 2021). Middle: We propose to use phases as a quantifiable diversity metric and generate datasets of trained models by varying their datasets $\mathcal{D}$, architectures $\mathcal{A}$ and training parameters $\lambda$ s.t. they cover all known phases across different domains. Right: Phases manifest beyond evaluation metrics in applications of pre-trained models, e.g., in weight averaging. Our Phase Model Zoos allow systematic evaluation of methods that use pre-trained models.
@@
To fill that gap, structured populations have been published as model zoo datasets, in domains as varied as computer vision (Unterthiner et al., 2020; Eilertsen et al., 2020; Schürholt et al., 2022b), adversarial robustness assessment (Croce et al., 2020), bioimaging (Ouyang et al., 2022) or Earth observation (Honegger et al., 2023). These populations, however, contain only models trained on one domain, and consider the diversity of the model in their populations only in terms of their generating factors. They report the distribution in model performance and model similarity, but it remains unclear whether that indicates conceptually different models, or whether they are variations of similar representations.
@@
  • We propose a new notion of diversity in model zoos by covering different model phases that relate to qualitatively different model behavior using loss landscape taxonomy.
@@ 2 Loss Landscape Taxonomy
The motivation for introducing phases and phase transitions in NN loss landscapes is rooted in statistical mechanics, where such phenomena explain qualitative changes in system behavior (Martin and Mahoney, 2019a). Phases represent distinct regions in the parameter space where the system’s properties are homogeneous or change smoothly, while phase transitions mark abrupt changes in these properties. In NNs, phases manifest in terms of generalization performance. A prominent example for such a phase transition is the double descent pheonmenon (Nakkiran et al., 2019), a phase transition along the axis of model size (Liao et al., 2020; Derezinski et al., 2020). Similar empirical observations have been made recently on the emergent behaviors of large language models (Wei et al., 2022), in which non-smooth transitions can occur when some training hyperparameters (such as the model size) are modified. However, it is not conclusive whether these emergent behaviors are indeed sharp phase transitions or merely due to specific ways of experimental measurements (Schaeffer et al., 2024). These phases and transitions are expected due to the complex, high-dimensional nature of NN optimization, where varying control parameters like data noise and training iterations can lead to qualitatively different behaviors, akin to physical systems undergoing phase changes. Motivated by statistical physics, Martin and Mahoney (2019a) identify two main types of hyperparameters in NN training: the noisiness of the training process, dubbed temperature, and the amount of data relative to the size of the model, dubbed relative load. Using that notion, distinct phases with qualitatively different model properties on the temperature-load landscape can be identified (Yang et al., 2021). Interestingly, the phases and phase transitions can be linked to the structure of the loss landscape (Yang et al., 2021). Specifically, metrics such as the training loss, the sharpness of local minima, and mode connectivity or prediction similarity computed on the training data can be used to identify the phase of a model. This, in turn, allows for inference of model quality and the design of training algorithms that adapt when the phases change (Zhou et al., 2023b).
@@
[…]
@@
  • Phase IV-A (generalizing): Train loss is small; Hessian metrics are relatively small; Mode connectivity is near-zero (no barrier between minima); CKA similarity is relatively large.
  • Phase IV-B (overfitted): Train loss is small; Hessian metrics are relatively small; Mode connectivity is near-zero; CKA similarity is relatively small.
@@ 3 Phase Transition Model Zoos
To create a population of models that covers relevant phases and can be used to evaluate for phase transitions, we train structured populations of NNs with several architectures on different datasets following the blueprint introduced by Unterthiner et al. (2020) and Schürholt et al. (2022b). Within each model zoo population, we systematically vary load-like and temperature-like hyperparameters to realize all of the phases. For every model in the zoo, our dataset includes multiple checkpoints (i.e. saved model weights), at different training epochs. We annotate these samples with performance metrics (training and test loss and accuracy), as well as the loss landscape metrics outlined in Section 2. We further track loss and accuracy on train and — if available — validation and test sets. In the following, we first detail our model zoo generation scheme, which we use to create 12 zoos on three domains: computer vision, natural language and physical systems. Further details can be found in Appendix A.2. Subsequently, we analyze our models with conventional performance metrics, but also with loss landscape metrics to quantify the qualitative diversity of the proposed zoos and validate that all of the phases are realized.
@@
[…]
@@
Lastly, we also create zoos of models for the simulation of physical systems. With that, we want to evaluate whether phases and phase transitions generalize even beyond the conventional deep-learning domains. Specifically, we train physics-informed neural networks (PINNs) (Raissi et al., 2019) to learn the solution to partial differential equations, following the implementation setup in (Krishnapriyan et al., 2021). To curate the dataset, we simulate the 1D convection partial differential equation (pde) and randomly sample the domain’s collocation points (position, time) as the data samples. Each data sample is comprised of position and time as input and the PDE solution value as the label. The training objective includes minimizing the loss of data prediction, boundary/initial conditions, and physics-based regularization. We train a 4-layer fully connected NN with 50 neurons per layer and tangent activation function using the L-BFGS optimizer. By varying temperature and load paramters, we train a zoo with 420 unique PINNs and 4200 checkpoints.
@@ 4.1 Phases Systematically Emerge Across all Zoos
A central reason we focus on load-like and temperature-like parameters is that they unify the effects of many hyperparameters into two axes. Empirically, models occupying the same phase on these axes exhibit qualitatively similar properties—regardless of which specific training factors (e.g., batch size, learning rate, or data size) gave rise to those load and temperature values. Consequently, rather than exhaustively exploring every hyperparameter combination, we can vary these two axes to capture all known phases of behavior. This is not only supported by prior work—where phase boundaries remain stable across different load–temperature choices ((Yang et al., 2021), Appendix D)—but also reflected in our results, suggesting strong generality of phase transitions to a broad range of architectures and training regimens.
@@
As illustrated in Figure 3 and in Figures 6-17, the phases manifest clearly in the combination of loss landscape metrics such as Hessian trace, mode connectivity, and CKA similarity. In particular, Phase IV-B, associated with the best test accuracy, is marked by low loss and high generalization performance. On the ResNet zoos, our results reveal that learning rate decay plays a significant role in shaping the phase distribution. Specifically, decaying the learning rate by $41e$ 4 under cosine annealing increases the area of Phase IV (well-trained regime) while reducing the presence of Phase II (under-trained regime), as the effect of batch size variations diminishes. This may be an indication of why learning rate decay is so successful. Our experiments show that the phase transitions generalize across different datasets, architectures, and training regimes. ViTs, for instance, display a sharper transition from Phase II to IV when trained without heavy augmentation (Figure 8). GPT-2 models exhibit a comparable pattern (Figure 16), suggesting that phase layouts remain consistent across distinct architectures, including transformers.
@@ 4.2 Phase Transitions in Neural Network Methods
Building on the observation that phase coverage describes performance and generalization during training, we now examine how phases influence four popular downstream methods: fine-tuning, pruning, ensembling, and weight averaging.
Figure 5: Phases and phase transitions of ResNet-18 models trained on CIFAR-10 in top: pre-training and loss-landscape metrics and bottom: downstream methods. Phase transitions emerge in downstream methods like fine-tuning, pruning and weight averaging.
In practice, many widely-used methods rely on pre-trained models—raising the question of whether and how a model’s phase continues to matter when its weights are reused. Below, we evaluate four canonical techniques and discuss how our phase-covering model zoos enable a more comprehensive analysis of their behavior. Across all these downstream applications, we find distinct phases in downstream performance, see Figure 5. Some of them overlap with the (pre-)training phases, some are distinct. These results are notable for three reasons: (i) it demonstrates that phase transitions exist broadly (ii) the load-temperature axis of our zoos is general enough to cover phase transitions in several downstream applications, beyond pre-training, significantly improving the applicability of our zoos; (iii) the datasets facilitate research of where these methods work or fail beyond point-wise evaluation, and build towards phase-aware method design.
@@
When fine-tuning, e.g. a CIFAR10-pretrained ResNet-18 on STL-10 (Figures 6,12, bottom left), we find that networks ending in a well-trained (Phase IV) regime yield stronger adaptation. Conversely, under-fitted or over-fitted phases struggle to reach comparable accuracy despite identical fine-tuning settings. This aligns with earlier observations that good phases not only benefit the original task but also confer more adaptable representations. For less aligned tasks like Tiny-Imagenet to STL-10 (Figure 13), we notice a phase shift from pre-trained to fine-tuned models. Understanding the relation between pre-training tasks and phase alignment using our model zoos can help pre-train or choose models to fine-tune in a more targeted fashion.
@@
Pruning has been shown to exhibit phase-dependent outcomes (Zhou et al., 2023b). We confirm that Phase IV networks often preserve accuracy better under high sparsity, while certain under-fitted/low-connectivity phases degrade quickly (Figure 5 bottom, second from left). By comparing across load-temperature grids, our model zoos help pinpoint where and why pruning works best.
@@
Combining model predictions (ensembling) has been investigated for phase transitions (Theisen et al., 2023). Our experiments confirm that ensembles tend to improve performance in load-temperature regions where local minima are well-connected (Figure 5 and 8). Notably, the phase of high-performance ensembles is larger than the individual model high-performance Phase IV, demonstrating the robustness benefit of ensembling with respect to pretraining settings. In practice, that corresponds to phases with near-zero mode connectivity barriers. Where mode connectivity is negative, ensembling provides less benefit. The zoo’s phase annotations thus indicate not just whether ensembling helps, but which configurations best exploit ensemble diversity.
@@
Averaging weights has recently gained popularity to improve robustness or combine knowledge (Izmailov et al., 2019; Wortsman et al., 2022a; Ilharco et al., 2022). We test averaging across training epochs (Wise-FT (Wortsman et al., 2022b)) and across pre-training seed (git re-basin (Ainsworth et al., 2022)). Our experiments (Figures 615) reveal distinct phase-dependent outcomes: Wise-FT is notably effective in high-temperature regions (Phases I or II) where averaging reduces noisy training updates, while in flatter phases (e.g., IV) it can yield little additional benefit. Git Re-Basin, on the other hand, hinges on positive mode connectivity; negative-connectivity phases (I or III) remain difficult to merge without performance loss. Hence, Hessian trace and mode connectivity strongly govern weight-averaging success, underscoring the value of our phase-annotated model zoos for pinpointing when and where averaging is most fruitful. Here in particular, point-evaluations of merging a few models lack the big picture necessary to build understanding.

### v2

@@ Abstract
Using the weights of trained Neural Network (NN) models as data modality has recently gained traction as a research field — dubbed Weight Space Learning (WSL). Multiple recent works propose WSL methods to analyze models, evaluate methods, or synthesize weights. Weight space learning methods require populations of trained models as datasets for development and evaluation. However, existing collections of models — called ‘model zoos’ — are unstructured or follow a rudimentary definition of diversity. In parallel, work rooted in statistical physics has identified phases and phase transitions in NN models. Models are homogeneous within the same phase but qualitatively differ from one phase to another. We combine the idea of ‘model zoos’ with phase information to create a controlled notion of diversity in populations. We introduce 12 large-scale zoos that systematically cover known phases and vary over model architecture, size, and datasets. These datasets cover different modalities, such as computer vision, natural language processing, and scientific ML. For every model, we compute loss landscape metrics and validate full coverage of the phases. With this dataset, we provide the community with a resource with a wide range of potential applications for WSL and beyond. Evidence suggests the loss landscape phase plays a role in applications such as model training, analysis, or sparsification. We demonstrate this in an exploratory study of the downstream methods like transfer learning or model weights averaging.
@@ 1 Introduction
Figure 1: Overview of the approach. Left: Phases of Neural Networks with qualitatively different properties have been identified (Martin and Mahoney, 2019a; Yang et al., 2021). Middle: We propose to use phases as a quantifiable diversity metric and generate datasets of trained models by varying their datasets $\mathcal{D}$, architectures $\mathcal{A}$ and training parameters $\lambda$ s.t. they cover all known phases across different domains. Right: Phases manifest beyond evaluation metrics in applications of pre-trained models, e.g., in weight averaging. Our Phase Model Zoos allow systematic evaluation of methods that use pre-trained models.
@@
There is growing research interest in addressing these gaps by analyzing collections of trained models. Using neural networks as data modality – dubbed weight space learning (WSL) has become its own research field [^1] ICLR 2025 hosts the first workshop on Neural Networks as Data Modality https://weight-space-learning.github.io/. Within the field of WSL, populations of models are used in a variety of domains: Applications include predicting how well models generalize (Jiang et al., 2020), or detecting poisoned models (IARPA, 2024). Other methods use WSL torwards Neural Architecture Search to identify the relation between weights and model performance (Martin et al., 2021; Zhou et al., 2024; Liu et al., 2024). Different work re-uses existing trained weights to extend conventional transfer learning strategies (Yosinski et al., 2014; Mensink et al., 2021) by combining multiple models, via a combination of their outputs (Polikar, 2012; Ganaie et al., 2022; Mohammed and Kora, 2023), knowledge distillation (Liu et al., 2019; Shu et al., 2021), or by directly combining their weights (Wortsman et al., 2022a, b; Mitchell et al., 2022; Ainsworth et al., 2022; Ilharco et al., 2022; Rame et al., 2023; Ramé et al., 2024; Navon et al., 2024), to build combined models with improved in- and out-of-distribution performance.
@@
Another line of research investigates phase transitions in NNs and connects them to loss landscape metrics (Martin and Mahoney, 2019a; Yang et al., 2021). Motivated by statistical physics, they identify two main types of hyperparameters in NN training: the noisiness of the training process, dubbed temperature, and the amount of data relative to the size of the model, dubbed relative load. Using that notion, they identify distinct phases with qualitatively different model properties on the temperature-load landscape (Yang et al., 2021). Phase transitions have been studied extensively in the machine learning literature (Schwarze et al., 1992; Seung et al., 1992; Martin and Mahoney, 2019a, b). While phase transitions are typically studied theoretically under certain limits, e.g., an infinitely wide NN (Lewkowycz et al., 2020), we observe qualitatively similar properties in models with a practical size. Within phases, models are relatively homogeneous, with abrupt changes from one phase to the next. Along the load dimension, one example of a phase transition is double descent (Nakkiran et al., 2019). Neural scaling laws describe the load and temperature relation, e.g., as power laws to train models to specific phases (Hestness et al., 2017; Rosenfeld et al., 2020; Gordon et al., 2021; Hoffmann et al., 2022; Sorscher et al., 2022).
@@
To fill that gap, structured populations have been published as model zoo datasets, in domains as varied as computer vision (Unterthiner et al., 2020; Eilertsen et al., 2020; Schürholt et al., 2022b; Falk et al., 2025a), adversarial robustness assessment (Croce et al., 2020), bioimaging (Ouyang et al., 2022) or Earth observation (Honegger et al., 2023b, a). These populations, however, contain only models trained on one domain, and consider the diversity of the model in their populations only in terms of their generating factors. They report the distribution in model performance and model similarity, but it remains unclear whether that indicates conceptually different models, or whether they are variations of similar representations.
@@
  • We propose a new notion of diversity in model zoos by covering different model phases that relate to qualitatively different model properties using loss landscape taxonomy.
@@
  • We make our dataset available at https://phasetransitions.modelzoos.cc
@@ 2 Loss Landscape Taxonomy
The motivation for introducing phases and phase transitions in NN loss landscapes is rooted in statistical mechanics, where such phenomena explain qualitative changes in system behavior (Martin and Mahoney, 2019a). Phases represent distinct regions in the parameter space where the system’s properties are homogeneous or change smoothly, while phase transitions mark abrupt changes in these properties. In NNs, phases manifest in terms of generalization performance. A prominent example for such a phase transition is the double descent pheonmenon (Nakkiran et al., 2019), a phase transition along the axis of model size (Belkin et al., 2019; Liao et al., 2020; Derezinski et al., 2020). Similar empirical observations have been made recently on the emergent abilities of large language models (Wei et al., 2022), in which non-smooth transitions can occur when some training hyperparameters (such as the model size) are modified. However, it is not conclusive whether these emergent abilities are indeed sharp phase transitions or merely due to specific ways of experimental measurements (Schaeffer et al., 2024). These phases and transitions are expected due to the complex, high-dimensional nature of NN optimization, where varying control parameters like data noise and training iterations can lead to qualitatively different properties, akin to physical systems undergoing phase changes. Motivated by statistical physics, Martin and Mahoney (2019a) identify two main types of hyperparameters in NN training: the noisiness of the training process, dubbed temperature, and the amount of data relative to the size of the model, dubbed relative load. Using that notion, distinct phases with qualitatively different model properties on the temperature-load landscape can be identified (Yang et al., 2021). Interestingly, the phases and phase transitions can be linked to the structure of the loss landscape (Yang et al., 2021), especially its global structure. We note that the study of global loss landscapes has been an active area of research, and it is generally understood that the properties of neural networks cannot be fully captured by local sharpness alone (Fort and Jastrzebski, 2019; Fort et al., 2020). Yang et al. (2021) present the first empirical attempt to quantify the transition from a globally well-connected to a globally less well-connected loss landscape. Specifically, metrics such as the training loss, the sharpness of local minima, and mode connectivity or representation similarity computed on the training data can be used to identify the phase of a model.
@@
where $t^{*}=\underset{t}{\mathrm{argmin}}\left|\frac{1}{2}(\mathcal{L}({\bm{\theta}})+\mathcal{L}({\bm{\theta}}^{\prime}))-\mathcal{L}(\gamma_{\phi}(t))\right|$. Here, $\texttt{mc}<0$ indicates a loss barrier between the two models and hence poor connectivity. $\texttt{mc}>0$ reveals lower loss regions between the models which indicates poor training. $\texttt{mc}\approx 0$ indicates well-connected models.
@@
  • Phase IV-A (overfitted): Train loss is small; Hessian metrics are relatively small; Mode connectivity is near-zero; CKA similarity is relatively small.
  • Phase IV-B (generalizing): Train loss is small; Hessian metrics are relatively small; Mode connectivity is near-zero (no barrier between minima); CKA similarity is relatively large.
Yang et al. (2021) use these metrics, along with fixed thresholds, to partition training into distinct phases. Training loss separates early phases (Phase I and II) from later ones (Phase III and IV). Mode connectivity, depending on whether it is positive or negative, differentiates Phase I from Phase II. Hessian trace and CKA similarity are then used to further divide Phase IV into two sub-phases. Zhou et al. (2024) formalize this procedure as a hierarchical decision tree, where thresholds for each metric can be learned using a small number of reference models. This optimization is performed using the bounded Brent method to maximize prediction accuracy on ‘failure modes’ of models. They show that the resulting phase boundaries can be transferred effectively to new tasks or model architectures.
@@ 3 Phase Transition Model Zoos
To create a population of models that covers relevant phases and can be used to evaluate for phase transitions, we train structured populations of NNs with several architectures on different datasets following the blueprint introduced by Unterthiner et al. (2020) and Schürholt et al. (2022b). Within each model zoo population, we systematically vary load-like and temperature-like hyperparameters to realize all of the phases. For every model in the zoo, our dataset includes multiple checkpoints (i.e. saved model weights), at different training epochs. We annotate these samples with performance metrics (training and test loss and accuracy), as well as the loss landscape metrics outlined in Section 2. Since the exact localization of phase boundaries is still a topic for ongoing research, we annotate each model with its loss landscape metrics as quantitative ground truth, from which different methods can derive slightly different phase transition boundaries. We further track loss and accuracy on train and — if available — validation and test sets. In the following, we first detail our model zoo generation scheme, which we use to create 12 zoos on three domains: computer vision, natural language and physical systems. Further details can be found in Appendix A.2. Subsequently, we analyze our models with conventional performance metrics, but also with loss landscape metrics to quantify the qualitative diversity of the proposed zoos and validate that all of the phases are realized.
@@
Ten computer vision zoos form the foundation of our dataset and demonstrate the generalization of the phase transitions accross architectures and datasets, while the language and SciML zoos show generalization of the phase transitions to other domains and tasks. We generate the computer vision zoos from combinations between two architectures $\{\textrm{ResNet},\textrm{ViT}\}$ of different sizes and four standard computer vision datasets $\{\textrm{SVHN},\textrm{CIFAR-10},\textrm{CIFAR-100},$ $\textrm{TinyImagenet}\}$ (Netzer et al., 2011; Krizhevsky and Hinton, 2009; Le and Yang, 2015). Details on the model zoo configurations can be found in Table 4 in the Appendix. We choose ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2021) architectures for the zoos because of their proliferation in computer vision to achieve representative populations. Importantly, ResNet and ViT architectures allow smooth scaling of model width and thus model capacity for the same architecture without the need to adjust the learning scheme. Similarly, the set of datasets evaluate generalization accross different data distributions. Taken together, the vision zoos contain $\sim$ 1.8k unique models and $\sim$ 55k checkpoints.
@@
Lastly, we also create zoos of models for the simulation of physical systems. With that, we want to evaluate whether phases and phase transitions generalize even beyond the conventional deep-learning domains. Specifically, we train physics-informed neural networks (PINNs) (Raissi et al., 2019) to learn the solution to partial differential equations, following the implementation setup in (Krishnapriyan et al., 2021). To curate the dataset, we simulate the 1D convection partial differential equation (pde) and randomly sample the domain’s collocation points (position, time) as the data samples. Each data sample is comprised of position and time as input and the PDE solution value as the label. The training objective includes minimizing the loss of data prediction, boundary/initial conditions, and physics-based regularization. We train a 4-layer fully connected NN with 50 neurons per layer and tangent activation function using the L-BFGS optimizer. By varying temperature and load paramters, we train a zoo with $700$ unique PINNs and 7 0007\,000 checkpoints.
@@ 4.1 Phases Systematically Emerge Across all Zoos
A central reason we focus on load-like and temperature-like parameters is that they unify the effects of many hyperparameters into two axes. Empirically, models occupying the same phase on these axes exhibit qualitatively similar properties—regardless of which specific training factors (e.g., batch size, learning rate, or data size) gave rise to those load and temperature values. Consequently, rather than exhaustively exploring every hyperparameter combination, we can vary these two axes to capture all known phases. This is not only supported by prior work—where phase boundaries remain stable across different load–temperature choices (Appendix D in Yang et al. (2021))—but also reflected in our results, suggesting strong generality of phase transitions to a broad range of architectures and training regimens.
@@
As illustrated in Figure 3 and in Figures 6-17, the phases manifest clearly in the combination of loss landscape metrics such as Hessian trace, mode connectivity, and CKA similarity. While different methods to locate phase transitions may identify slightly differ boundaries, the evaluation of the performance and loss landscape metrics in the phase plots clearly demonstrates the coverage of the phases described in previous work. In particular, Phase IV-B, associated with the best test accuracy, is marked by low loss and high generalization performance. On the ResNet zoos, our results reveal that learning rate decay plays a significant role in shaping the phase distribution. Specifically, decaying the learning rate by $1e4$ under cosine annealing increases the area of Phase IV (well-trained regime) while reducing the presence of Phase II (under-trained regime), as the effect of batch size variations diminishes. This may be an indication of why learning rate decay is so successful. Our experiments show that the phase transitions generalize across different datasets, architectures, and training regimes. ViTs, for instance, display a sharper transition from Phase II to IV when trained without heavy augmentation (Figure 8). GPT-2 models exhibit a comparable pattern (Figure 16), suggesting that phase layouts remain consistent across distinct architectures, including transformers.
@@ 4.2 Phase Transitions in Neural Network Methods
Building on the observation that phase coverage describes performance and generalization during training, we now examine how phases influence four popular downstream methods: fine-tuning, pruning, ensembling, and weight averaging. In practice, many widely‐used methods rely on pre‐trained models—raising the question of whether and how a model’s phase continues to matter when its weights are reused. Below, we evaluate four canonical techniques and discuss how our phase‐covering model zoos enable a more comprehensive analysis of their properties. Across all these downstream applications, we find distinct phases in downstream performance, see Figure 5. Some of them overlap with the (pre-)training phases, some are distinct. These results are notable for three reasons: (i) it demonstrates that phase transitions exist broadly (ii) the load-temperature axis of our zoos is general enough to cover phase transitions in several downstream applications, beyond pre-training, significantly improving the applicability of our zoos; (iii) the datasets facilitate research of where these methods work or fail beyond point-wise evaluation, and build towards phase‐aware method design.
Figure 5: Phases and phase transitions of ResNet-18 models trained on CIFAR-10 in top: pre-training and loss-landscape metrics and bottom: downstream methods. Phase transitions emerge in downstream methods like fine-tuning, pruning and weight averaging. We use a different color palette for the downstream methods figures in order to make them easily identifiable.
@@
When fine‐tuning, e.g. a CIFAR10‐pretrained ResNet-18 on STL‐10 (Figures 6,12, bottom left), we find that networks ending in a well‐trained (Phase IV) regime yield stronger adaptation. Conversely, under‐fitted or over‐fitted phases struggle to reach comparable accuracy despite identical fine‐tuning settings. This aligns with earlier observations that good phases not only benefit the original task but also confer more adaptable representations. For less aligned tasks like Tiny-Imagenet to STL-10 (Figure 13), we notice a phase shift from pre-trained to fine-tuned models. Understanding the relation between pre-training tasks and phase alignment using our model zoos can help pre-train or choose models to fine-tune in a more targeted fashion.
@@
Pruning has been shown to exhibit phase‐dependent outcomes (Zhou et al., 2023b). We confirm that Phase IV networks often preserve accuracy better under high sparsity, while certain under‐fitted/low‐connectivity phases degrade quickly (Figure 5 bottom, second from left). By comparing across load‐temperature grids, our model zoos help pinpoint where and why pruning works best.
@@
Combining model predictions (ensembling) has been investigated for phase transitions (Theisen et al., 2023). Our experiments confirm that ensembles tend to improve performance in load‐temperature regions where local minima are well‐connected (Figure 5 and 8). Notably, the phase of high-performance ensembles is larger than the individual model high-performance Phase IV, demonstrating the robustness benefit of ensembling with respect to pretraining settings. In practice, that corresponds to phases with near‐zero mode connectivity barriers. Where mode connectivity is negative, ensembling provides less benefit. The zoo’s phase annotations thus indicate not just whether ensembling helps, but which configurations best exploit ensemble diversity.
@@
Averaging weights has recently gained popularity to improve robustness or combine knowledge (Izmailov et al., 2019; Wortsman et al., 2022a; Ilharco et al., 2022). We test averaging across training epochs (Wise-FT (Wortsman et al., 2022b)) and across pre-training seed (git re-basin (Ainsworth et al., 2022)). Our experiments (Figures 615) reveal distinct phase‐dependent outcomes: Wise‐FT is notably effective in high‐temperature regions (Phases I or II) where averaging reduces noisy training updates, while in flatter phases (e.g., IV) it can yield little additional benefit. Git Re‐Basin, on the other hand, hinges on positive mode connectivity; negative‐connectivity phases (I or III) remain difficult to merge without performance loss. Hence, Hessian trace and mode connectivity strongly govern weight‐averaging success, underscoring the value of our phase‐annotated model zoos for pinpointing when and where averaging is most fruitful. Here in particular, point-evaluations of merging a few models lack the big picture necessary to build understanding.